OpenStudy (anonymous):

what is the slope of a line that is perpendicular to the line represented by the equation y-5x=5? 5 1/5 -1/5 -5

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

and why did u become my fan

OpenStudy (anonymous):

thanks, because you answered

OpenStudy (anonymous):

except -5 was wrong

OpenStudy (anonymous):

actually -5 is wrong

OpenStudy (anonymous):

sorry i thought it was parralell

OpenStudy (angela210793):

two lines are perpendicular when their slopes have the product=-1 so y-5x=5-->y=5+5x so the slope of the line perpendicular with the line given is 5*x=-1 x=-1/5

OpenStudy (angela210793):

so the slope must be =-1/5

OpenStudy (anonymous):

slope of \[y=5x+5\] is 5 slope of perp line is \[-\frac{1}{5}\]

OpenStudy (anonymous):

thanks, I'll do that

OpenStudy (anonymous):

the given line y-5x=5 can also be represented as y=5x+5. comparing this equation with y=mx+c,where m is the slope,we get the slope as 5. the condition for two lines to be perpendicular is,product of slopes should be -1. so the slope of the line perpendicular to the gin=ven equation would be,(-1/5).